Please help me with the problem below:
Grandma was riding her Harley through the desert when she remembered that she left her knitting needles at the bar. The road runs 15 miles west and 30 miles south of the bar where Grandma is currently. If she can ride 50 miles per hour on road and 30 miles per hour on sand, and she begins by first backtracking up the road, what route will get her back the quickest?

Answer by rothauserc(4718)

Backtracking, she has 45 miles of road which she can cover using rate*time = distance,
50 * time = 45
Time = 45/50 = 9/10 of an hour
Now, instead of backtracking, she heads out over the sand
D^2 = 15^2 + 30^2
D^2 = 225 + 900 = 1125
D = square root (1125) = 33.54
30 * time = 33.54
Time = 33.54 / 30 = 1.12 hours
Backtracking along the road will get her back sooner